a car is going on a national highway with a constant accerleration .it crosses certain milestone with a speed of 7m/s ..it crosses the next mile stone with a speed of 17m/s. with what speed it crosses the next milestone?
please answer it in elloberately so that i can understand it :)

Grade:11

1 Answers

Arjun Sharma
6 years ago
Dist. Covered =1600mt.Lets consider second case when car covers 2nd milestone with speed of 17m/s.initial velocity = 7m/sFinal velocity= 17m/sSince 2as=v^2 -u^22a*1600=17^2-7^2So acceleration=3/40 unitsNow lets consider 3rd milestoneWhere u=17 a=3/40unitsAgain 2as+u^2=v^22*3/40*1600+17^2=v^2After calculation we shall getv^2=529So v= 23m/sHence ans. Is 23 m/s